# VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

### v1

@@ 4 Experiments
4.1 Setup
In this section, we detail our experimental configuration, including experimental datasets, model specifications, and baseline methods for evaluating the performance of the proposed VisPath framework in generating visualization code from natural language queries.
4.1.1 Experimental Datasets
We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al. (2024) and the Qwen-Agent Code Interpreter Benchmark. MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries. In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
4.1.2 Models Used
Large Language Models (LLMs):
For the code inference stage, we experiment with GPT-4o mini Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) to generate candidate visualization code from the reasoning paths. Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with recommendations from previous work Yang et al. (2024). To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) as our visualization feedback model, which provides high-quality reference assessments.
Vision Language Models (VLMs):
4.1.3 Evaluation Metrics
In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability. MatPlotBench assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions. Qwen-Agent Code Interpreter benchmark evaluates visualization models based on Visualization-Hard and Visualization-Easy, measuring how well generated images align with queries of different difficulty levels. Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric. Previous studies show GPT-based VLM evaluations align with human assessments, hence VLM was used for evaluation.
@@ 4.1.4 Baseline Methods
We compare VisPath against competitive baselines. Specifically, (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses chain-of-thought prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al. (2024) first expands the query and then refines the code via a self-debugging loop with feedback. For a fair comparison, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well, and VisPath generates three reasoning paths with corresponding visual feedback to refine the final output. [^2] Prompts are detailed in Appendix A.
@@ 4.2 Experimental Analysis
In this section, we present a detailed evaluation of our proposed VisPath framework, summarizing the performance results against four baselines: (1) Zero-Shot, (2) Chain-of-Thought (CoT) Prompting, (3) Chat2VIS, and (4) MatPlotAgent.
From Table  1, we find that the Zero-Shot method, which directly generates visualization code from the user query without intermediate reasoning, suffers from ambiguous interpretations and error-prone outputs. These issues are particularly pronounced when the input queries are underspecified, contain implicit assumptions, or require complex reasoning steps. CoT Prompting, on the other hand, mitigates some of these challenges by employing chain-of-thought (CoT) reasoning, which helps the model articulate its decision-making process step by step. While this structured reasoning approach improves interpretability and correctness, its reliance on a single reasoning trajectory limits its capacity to explore alternative solutions.
Meanwhile, Chat2VIS builds upon CoT Prompting by incorporating guided prefixes to better structure user queries and disambiguate input intent. This enhancement leads to more coherent code generation and reduces errors stemming from unclear specifications. However, its effectiveness still depends on the predefined guiding templates, which may not fully adapt to highly variable or novel query structures. MatPlotAgent further refines this approach through query expansion and an iterative self-debugging loop that enhances code robustness. However, its reliance on an iterative correction mechanism comes at the cost of computational efficiency and does not fully leverage diverse reasoning strategies that could lead to more creative and contextually accurate solutions.
In contrast, our novel framework, VisPath, overcomes these challenges by dynamically generating multiple reasoning paths. By exploring diverse interpretations of user intent simultaneously, VisPath significantly enhances the robustness of the generated visualization code. Furthermore, it integrates structured feedback from Vision-Language Models (VLMs) to refine its outputs, ensuring both higher execution accuracy and improved aesthetic quality. This feedback-driven optimization resolves ambiguities more effectively and selects the most appropriate solution from a wide range of possibilities. Evaluations conducted on benchmark datasets such as MatPlotBench and Qwen-Agent Code Interpreter benchmark demonstrate that VisPath consistently outperforms the baseline methods by average 17%. VisPath exhibits a superior ability to handle complex and ambiguous visualization requests by synthesizing multiple reasoning trajectories and incorporating structured feedback, making it a highly effective tool for automated visualization generation.
@@ 4.3 Ablation Study
This section further explores the impact of varying the number of reasoning paths on the performance of visualization code generation. We conduct a series of ablation studies to demonstrate the robustness of our VisPath framework under alternative settings. Specifically, we investigate (i) the effect of varying the number of generated reasoning paths and (ii) the impact of a simplified integration strategy for synthesizing the final visualization code.
@@ 4.3.1 Varying the Number of Reasoning Paths
To further assess the contribution of reasoning path diversity, we conducted ablation experiments varying the number of generated paths $K\in\{2,3,4\}$ to evaluate their effects on execution accuracy, visualization quality and interpretability. Evaluation has been conducted on both the MatPlotBench and Qwen-Agent Code Interpreter benchmark datasets. The study reveals an interesting pattern observed when the number of reasoning path is reduced. As illustrated in Table 2 and Table 3, when $K=2$, the system exhibited limited diversity in which occasionally results in missing nuanced interpretations of complex user queries.
Model K MatPlotBench
Plot Score Executable Rate (%)
GPT-4o mini 2 64.02 58
3 66.12 60
4 64.68 62
Gemini 2.0 Flash 2 56.59 56
3 59.37 61
4 54.60 54
Table 2: Results of ablation study on MatPlotBench.
Model K Qwen-Agent
Hard Easy Avg.
GPT-4o mini 2 67.50 55.72 61.61
3 70.68 57.23 63.96
4 66.42 50.25 58.34
Gemini 2.0 Flash 2 74.21 57.50 65.86
3 80.79 57.17 68.98
4 76.41 54.50 65.46
Table 3: Results of ablation study on Qwen-Agent Code Interpreter benchmark.
Conversely, increasing $K$ to 4 introduces extra paths which can sometimes add noise as less relevant interpretations are considered. Specifically, we observed cases where redundant or overly complex visualization components were generated, hence making the final output harder to be interpreted. While the additional reasoning paths increased the overall exploration space, they also required more extensive filtering and selection, leading to suboptimal execution efficiency.
Notably, $K=3$ consistently achieves the best balance by providing sufficient diversity without introducing excessive noise. With too few reasoning paths, some nuanced aspects of ambiguous queries are lost, while too many paths configuration result in unnecessary exploration of low relevancy solutions. The number of reasoning path diversity set as $K=3$ optimally provides sufficient diversity without overcomplicating the reasoning process.
4.3.2 Robustness with a Simple Integration Strategy
To further validate the robustness of VisPath, we also evaluate an alternative integration strategy that simplicies the aggregation of multiple reasoning paths. In contrast to our full feedback-driven iterative approach which refines visualization code through multiple rounds of vision-language feedback, this streamlined method directly aggregates candidate outputs from different reasoning trajectories without intermediate corrections.
Specifically, in this alternative strategy, three candidate codes, each derived from different reasoning paths, are directly combined to generate the final visualization output. This methodology reduces computational overhead and processing time while maintaining a degree of interpretability and correctness. Through examining the impact of this approach, we gain insight into the extent to which VisPath’s core strength stems from its multi-path reasoning capability versus its iterative refinement process.
@@ 4.3.2 Robustness with a Simple Integration Strategy
Table 4: Performance comparison of VisPath with and without visual feedback. The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two llm models.
As shown in Table 4, even under this simplified integration, the performance of VisPath significantly outperforms all baseline methods, demonstrating its robustness and adaptability. This result suggests that the diversity of reasoning paths alone ensures a resilient and effective outcome, even in the absence of extensive intermediate corrections. This finding shows that the strength of our framework largely stems from its multi-path reasoning design. Even when the integration is simplified, the diversity of reasoning paths ensures that the aggregated output remains robust and effective, thus showing the overall adaptability and effectiveness of our approach.
@@ 5 Discussion
Our main experiment and ablation studies decisively demonstrate that VisPath’s multi-path exploration revolutionizes visualization code generation by making it not just robust, but truly interpretable.  [^3] Case study is provided in Appendix B. Unlike traditional methods that often yield simplistic, one-dimensional results, VisPath embraces complexity—delving deeper to capture nuances of user intent that would otherwise be lost. By integrating multiple perspectives during reasoning, our framework ensures that final visualizations are not only data-accurate but also rich in context, inherently intuitive, and unmistakably aligned with the user’s true objectives.
The real breakthrough with VisPath is how it transforms visualization into a structured yet fluid process. Traditional methods often act like static translators, converting user queries into plots while neglecting the vital auxiliary components—legends, labels, line styles, and other elements—that enable true interpretability. They assume that a single pass can sufficiently capture the essence of a visualization, but this leads to visualizations that might technically be correct but fail at communication. In contrast, VisPath’s multi-path approach systematically explores different ways to represent data, integrating feedback at each step. This ensures that even subtle yet critical design elements are not just included, but optimized, making the resulting visualizations feel complete, polished, and immediately comprehensible.
Beyond accuracy, VisPath introduces a level of adaptability that is indispensable in today’s fast-evolving data landscape. It’s not just a tool for static reports—it is built for real-time streaming data, dynamic dashboards, and interactive analytics. Even with broad or loosely defined prompts, our framework intelligently synthesizes high-quality visualizations, closing the gap between intent and execution. As industries continue to demand clearer, faster, and more adaptable insights, VisPath is poised to lead the way, setting a new standard for intelligent, context-aware data visualization.
@@ 6 Conclusion
In this work, we present VisPath, a groundbreaking framework that redefines the landscape of automated visualization code generation over existing methods. Unlike prior methods, our approach seaminglessly combines multi-path reasoning with feedback-driven optimization. Accurately capturing diverse user intents and refine generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark. By prioritizing adapability and user intent alignment, VisPath is uniquely positioned to handle compledities of real-world data visualization tasks. Looking ahead, future work could explore VisPath’s adaptability in more dynamic, real-world visualization scenarios, further broadening its scope and enhancing its practical utility in complex data analysis contexts.
@@ 7 Limitations
Despite its effectiveness, the current iteration of our framework relies on a limited feedback mechanism, evaluating visualizations primarily based on two aspects: query-code alignment and query-plot image alignment. While these aspects provide valuable insights, they may not fully capture the granular elements that contribute to overall interpretability. To enhance the depth of feedback itself, future research could focus on decomposition of assessment process, assessing individual plot components separately. This granular analysis would render for a more nuanced assessment of readability, element appropriateness and visual coherence, ultimately leading to more refined visualization code generation. Strengthening the feedback mechanism in such way will be crucial for maximizing VisPath’s effectiveness in diverse and complex visualization scenarios.
@@
Prompt for Multi-Path Reasoning [System Prompt] According to the user query, expand and solidify the query into detailed instruction on how to write python code to fulfill the user query’s requirements. Import the appropriate libraries. Pinpoint the correct library functions to call and set each parameter in every function call accordingly. [User Prompt] Think step by step. Generate three distinct extended queries based on the given query. Ensure that you first analyze the given data description and create queries that align with the data. If no data description is provided, follow the original query as is. You must follow the Python list format for the output. Do not modify the detailed instructions from the original user query. Original query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
@@
Figure 3: Overview of end-to-end process of visualizing women’s million-dollar earnings data using VisPath. It includes the user query, raw data, multi-path reasoning, generated figures, visual feedback, final script, and the completed visualization.

### v2

@@ 4.1.4 Baseline Methods
We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al. (2024) first expands the query and then refines the code via a self-debugging loop with feedback. Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output. [^3] Prompts are detailed in Appendix A. For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well.
@@ 4.2 Experimental Analysis
VisPath is evaluated against four baselines: Zero Shot prompting, CoT prompting, Chat2VIS, and MatPlotAgent. The evaluation is conducted on MatPlotBench and the Qwen-Agent Code Interpreter benchmark using GPT-4o mini and Gemini 2.0 Flash, as shown in Table 1.
Zero-Shot prompting generates visualization code directly from natural language queries without intermediate reasoning. While computationally efficient, it often struggles to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs. On MatPlotBench (GPT-4o mini), it achieves a Plot Score of 62.38 and an Executable Rate of 53%. CoT prompting further introduces a single reasoning step to expose intermediate decisions and improve interpretability. However, on MatPlotBench, it slightly underperforms Zero-Shot in both Plot Score and Executable Rate, indicating reliance on a fixed reasoning path may reduce adaptability to diverse input structures.
Chat2VIS extends CoT prompting by adopting prefix templates to improve coherence and reduce ambiguity in user instructions. While this approach is effective for well-structured or common query formats, its dependence on fixed templates limits adaptability when processing loosely specified or novel queries. Such limitation is evident in its performance on MatPlotBench, where it achieves a Plot Score of 56.98 and an Executable Rate of 53%. Furthermore, MatPlotAgent incorporates query expansion and iterative self-debugging mechanisms to enhance robustness. While effective at correcting execution-level errors, its revisions are confined to localized adjustments and do not address higher-order semantic ambiguities.
In contrast, our proposed framework VisPath is specifically designed to overcome these limitations observed in prior methods by dynamically generating multiple reasoning paths and refining them through structured visual feedback. In particular, template-based approaches such as Chat2VIS offer limited adaptability due to their reliance on predefined input formats, while methods such as MatPlotAgent focus on localized corrections without addressing broader semantic ambiguity. Unlike prior methods, VisPath generates diverse interpretations of user intent and evaluates them holistically using structured vision-language feedback. This enables more flexible handling of under-specified or ambiguous inputs, resulting in semantically aligned and executable visualizations.
Evaluated across multiple benchmark settings, VisPath notably outperforms baselines, achieving up to 9.14 point gains in Plot Score and a 10% point increase in Executable Rate. These improvements well demonstrate VisPath’s robustness in exploring diverse reasoning paths and iteratively refine outputs through structured visual feedback, effectively reducing semantic ambiguity and improving execution reliability.
@@ 4.3 Ablation Study
To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments in this section. Specifically, we analyze the following three aspects: (i) varying the number of generated reasoning paths, (ii) the effect of removing visual feedback during integration, and (iii) the contribution of visual feedback beyond binary executability.
Figure 3: Effect of varying the number of reasoning paths $K$ on performance across datasets and models. Metrics include Plot Score, Executable Rate. The results show that $K=3$ achieves the best overall balance, with larger $K$ values reducing performance.
@@ 4.3.1 Varying the Number of Reasoning Paths
To investigate the contribution of reasoning path diversity, we conducted ablation experiments by varying the number of generated reasoning paths $K$. In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path multiplicity on the overall performance of VisPath, as shown in Figure 3.
We observe a consistent pattern across all model and dataset combinations: performance improves as $K$ increases from 2 to 3, confirming that limited diversity $(K=2)$ often fails to capture nuanced interpretations of user queries. For example, with GPT-4o mini on MatPlotBench, the Plot Score improves from 64.02 to 66.12 (+2.10), and the Executable Rate improves from 58% to 60% (+2 points). On Gemini 2.0 Flash, the Plot Score increases from 56.59 to 59.37 (+2.78), and the Executable Rate from 56% to 61% (+5 points).
While $K=4$ achieves the highest executable rate on MatPlotBench with GPT-4o mini (62%), we further extend our analysis up to $K$ = 8 to comprehensively assess the impact of reasoning path diversity. However, beyond $K=4$, we observe diminishing returns and even performance degradation, which is likely due to noisy or redundant reasoning paths. While added diversity initially aids interpretation, excessive expansion burdens the integration process and reduces overall efficiency.
Among all configurations tested up to $K=8$, $K=3$ emerges as the most balanced choice, offering substantial performance gains in both the Executable Rate and the Plot Score while avoiding the inefficiencies observed at higher values of $K$. Hence, we adopt $K=3$ as the default configuration throughout our experiments.
4.3.2 Robustness with a Simple Integration
We evaluate an alternative integration strategy that simplifies the aggregation of multiple reasoning paths to further validate the robustness of VisPath.
@@ 4.3.2 Robustness with a Simple Integration
Table 2: Performance comparison of VisPath with and without visual feedback. The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two LLMs.
Instead of refining each candidate visualization with feedback from vision-language feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections. This setup reduces computational overhead and execution time while preserving the benefits of interpretive diversity.
As shown in Table 2, even under this simplified configuration, VisPath outperforms all baseline methods, confirming that Multi-Path Reasoning alone offers a strong foundation for visualization code generation. While full feedback-driven optimization leads to additional performance improvements, this result highlights that the primary strength of VisPath lies in its capacity to explore and leverage diverse reasoning trajectories. The framework remains effective and adaptable, even with minimal refinements, further validating importance of its core design.
4.3.3 Distinct Contribution of Visual Feedback
To assess the role of visual feedback in improving code quality, we compare two variants of our framework. The first, VisPath (w/ feedback), uses a VLM to evaluate both rendered plots and error messages. The second, VisPath Execute (w/ binary feedback), simplifies evaluation by relying solely on the binary success or failure of code execution.
Incorporating rendered visual feedback improves the Executable Rate by $2\%-4\%$ and consistently boosts the Plot Score across both LLMs, as shown in Table 3. On GPT-4o mini, Plot Score increases from 64.82 to 66.12 (+1.30) and Executable Rate from 58% to 60% (+2 points). On Gemini 2.0 Flash, Plot Score rises from 57.68 to 59.37 (+1.69), and Executable Rate from 59% to 63% (+4 points). Despite the numerical gains are modest, the results demonstrate the unique value of structured visual evaluation. Visual feedback enables more refined and user-aligned outputs by capturing subtle rendering issues that may not affect executability, demonstrating its importance in the final synthesis stage.
LLM Variant Plot Score Executable Rate (%)
GPT-4o mini (w) feedback 66.12 60
(w) binary feedback 64.82 58
Gemini 2.0 Flash (w) feedback 59.37 63
(w) binary feedback 57.68 59
Table 3: Ablation results isolating the impact of rendered visual feedback. Comparing the full VisPath model using structured plot-based feedback with a binary-only feedback variant. Visual feedback provides consistent gains in both Plot Score and Executable Rate across LLMs.
@@ 5 Discussion
Our proposed VisPath framework substantially advances visualization code generation by addressing the core weaknesses of existing methods: limited interpretive flexibility and insufficient refinement. By employing Multi-Path Reasoning, VisPath explores diverse interpretations of user intent, which leads to more accurate visualizations, especially for ambiguous queries. Experimental results confirm its superiority: VisPath outperforms all baselines on both MatPlotBench and the Qwen-Agent Code Interpreter benchmark, with up to 9.14% improvement in Plot Score and 10% in Executable Rate. Ablation studies further validate VisPath’s design. First, increasing the number of reasoning paths enhances both visual quality and code executability. Second, even without visual feedback, Multi-Path Reasoning alone proves highly effective. Third, using structured plot-based feedback, rather than binary execution signals, significantly improves output alignment with user intent, confirming the value of our feedback-driven optimization module.
The figures provided in Appendix clearly illustrate VisPath’s robustness. [^4] The detailed cases are provided in Appendix B. In Case 1, it correctly centers conditional means using adjusted strip widths, unlike other methods that misalign pointplots or ignore key parameters. In Case 2, VisPath generates a proper polar bar chart with radial coordinates and legible labels, while others either fail to render or produce standard rectangular plots, with text often overlapping or unreadable. Case 3 highlights VisPath’s ability to handle compositional visualization requests: it correctly interprets the instruction to "visualize in 3 different ways" by generating a multi-subplot layout that includes line, scatter, and bar plots. Other approaches, by contrast, either collapsed all time series into a single chaotic plot or failed to differentiate the visualization modalities at all. These examples demonstrate how VisPath’s Multi-Path Reasoning and visual feedback lead to more precise and semantically aligned visualizations than current alternatives.
@@ 6 Conclusion
In this work, we present VisPath, a framework that leverages Multi-Path Reasoning and feedback-driven optimization to enhance automated visualization code generation. Unlike prior methods, our approach seamlessly combines Multi-Path Reasoning with feedback-driven optimization. Accurately capturing diverse user intents and refining generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark. By prioritizing adaptability, VisPath is uniquely positioned to handle ambiguous user queries through a combination of diverse reasoning paths and visual feedback integration. Future work could explore VisPath’s adaptability in more dynamic, real-world scenarios, further broadening its scope and practical utility in complex data analysis contexts.
@@ 7 Limitations
Despite its effectiveness, the current framework relies on a limited feedback mechanism focused on query-code and query-plot alignment. While informative, these signals may overlook finer-grained elements essential to interpretability. Thus, future work could improve feedback depth by assessing individual plot components, such as readability and visual coherence, enabling more precise and refined visualization code generation. Moreover, while achieving strong performance, VisPath requires several rounds of agent interaction, including multi-path reasoning, execution, and feedback integration, which may introduce inefficiencies in certain use cases. Future work could explore ways to selectively identify the most promising reasoning paths early in the process, reducing redundant computation while preserving the benefits of diverse interpretation.
@@
Prompt for Multi-Path Reasoning [System Prompt] According to the user query, expand and solidify the query into detailed instruction on how to write python code to fulfill the user query’s requirements. Import the appropriate libraries. Pinpoint the correct library functions to call and set each parameter in every function call accordingly. [User Prompt] Think step by step. Generate three distinct extended queries based on the given query. Ensure that you first analyze the given data description and create queries that align with the data. If no data description is provided, follow the initial query as is. You must follow the Python list format for the output. Do not modify the detailed instructions from the original user query. initial query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
@@
Figure 4: Overview of the end-to-end process of visualizing conditional means alignment.
Figure 5: Overview of the end-to-end process of generating a polar bar chart.
Figure 6: Overview of the end-to-end process of visualizing time series data in three different ways.